# Algorithms for Optimizing Acyclic Queries

### v2

@@ 6 $e_{2}=(l,r_{2}\leftarrow\mathsf{LA}(r_{j},d));w_{2}=\omega(e_{2})$
Starting from an $\alpha$-acyclic query $\mathcal{Q}$, Section˜C.1 describes how to convert it into a hypergraph $H$. Section˜C.2 describes how to convert it into a weighted line graph $L(H)$. Both conversions take linear time $\mathcal{O}(|\mathcal{Q}|)$ with respect to the query size as defined in Appendix˜C. Algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)\leq\mathcal{O}(|\mathcal{Q}|)$. With the inputs $L(H)$ and $T$, we can run Algorithm˜3 to construct $G^{\equiv}(H^{*})$ in $\mathcal{O}(m+n)\leq\mathcal{O}(|\mathcal{Q}|)$ (where $n$ is the number of vertices and $m$ is the number of edges in $L(H)$). Finally, we apply the spanning tree enumeration algorithm by Kapoor and Ramesh [22] on $G^{\equiv}(H^{*})$ to enumerate all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)\leq\mathcal{O}(|\mathcal{Q}|+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
4.3 $\gamma$-Acyclic Queries
The run time of Algorithm˜3 depends on the size of the line graph which can be quadratically larger than the input hypergraph. But if $H$ is $\gamma$-acyclic, we can bring the total time complexity of enumeration down to $\mathcal{O}(|H|+|\mathcal{T}(H)|)$. The line graph of a $\gamma$-acyclic hypergraph $H$ is its union join graph [25] $L(H)=\mathcal{U}(H)$, i.e., every edge $e\in E(L(H))$ in the line graph is an MST edge. By modifying the MCS algorithm to Algorithm˜7 in Appendix˜F, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$, namely the union join graph, in $\mathcal{O}(|H|+|E(\mathcal{U}(H))|)$. We simplify Algorithm˜3 to Algorithm˜8 in Appendix˜F, which compares the LCA edges $\lambda(e)$ of each non-tree edge $e$ and slides it to construct the equivalent graph $G^{\equiv}(H)$ in $\mathcal{O}(|E(\mathcal{U}(H))|)$. As $\gamma$-acyclicity guarantees that every non-tree edge $e$ is an MST edge, we can use the weights of the LCA edges to determine how to slide $e$. Let $\lambda(e)=\{e_{l},e_{r}\}$. If $e_{l}=e_{r}$, we slide $e$ to be parallel to $e_{l}$. Otherwise, if $e_{l}$ and $e_{r}$ have different weights, we slide $e$ to be parallel to the heavier one. If they have the same weight, we slide $e$ to form a triangle with them.
The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(\mathcal{U}(H))|-|E(T)|)$. The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(\mathcal{U}(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
5 The Canonical Join Tree of a Berge-Acyclic Query
An acyclic query can have exponentially many join trees with respect to its size. For example, the line graph of a clique query with $n$ relations is an $n$-clique $K_{n}$ with $n^{n-2}$ join trees by Cayley’s formula [7]. Enumerating all join trees can be prohibitive for large queries. On the other hand, the query optimizer does not need to consider all possible join trees to achieve good performance. For example, the implementation of Yannakakis’ algorithm by Zhao et al. [47] has similar performance on any join tree rooted at the largest relation. An alternative to enumeration is therefore to simply construct an arbitrary join tree for a given root. This can be done in linear time by the Maximum Cardinality Search (MCS) algorithm from a chosen relation as shown in Algorithm˜1. In this section, we show that the MCS algorithm produces a shallowest tree for a Berge-acyclic query, in the sense that the depth of every node in the tree is minimized. We prove that this shallowest tree is unique, and therefore call it the canonical join tree:
Definition 25.
A shallow join tree has practical benefits. For example, the depth of the join tree determines the number of sequential steps required in a parallel join algorithm. A shallow join tree tends to be wide and have more leaves, allowing better utilization of indices.
Although Berge-acyclicity was thought to be too restrictive when it was first introduced to database theory [13], we found it to be general enough to cover almost all acyclic queries encountered in the wild. As shown in Table˜1 of Appendix˜F, among 10 45410\,454 queries from five popular benchmarks, $9285$ are $\alpha$-acyclic, and only 8 of these are not Berge-acyclic. In retrospect, this should not be surprising, as most joins in relational databases are over primary/foreign keys. Emerging workloads in graph databases usually involve simple graphs and seldom require composite key joins. A query without composite key joins admits a linear hypergraph, where each pair of hyperedges shares at most one vertex. The following result establishes an equivalence between $\alpha$-acyclicity with linearity and Berge-acyclicity.
Proposition 26.
By Section˜5, every edge in the line graph of a Berge-acyclic hypergraph is labeled with a single variable, thus having a weight of 1. Every spanning tree is a maximum spanning tree, therefore a join tree.
Corollary 27.
5.1 Existence & Uniqueness of the Canonical Join Tree
The key insight leading to the existence and uniqueness of the canonical join tree is that the line graph of a Berge-acyclic hypergraph is geodetic [31], meaning that there is a unique shortest path between any pair of vertices. To prove this, we first review some relevant concepts and properties of graphs. All graphs below are simple and undirected.
Definition 28.
A cycle in a graph is a sequence of distinct vertices $v_{0},\ldots,v_{k-1\geq 2}$ such that there is an edge between $v_{i}$ and $v_{(i+1)\bmod k}$ for all $1\leq i\leq k$.
A special class of graphs called chordal graphs are intimately related to acyclic hypergraphs:
Definition 29.
A graph is chordal if every cycle of length at least 4 has a chord, i.e., an edge that is not part of the cycle but connects two vertices of the cycle.
The key step in this section is to show the line graph of a Berge-acyclic hypergraph is a block graph, which is a special class of chordal graphs. There are many different characterizations of block graphs [18], and we present a recent one:
Definition 30 (Block Graph [10]).
A graph $G=(R,E)$ with the vertex set $R$ and the edge set $E$ is a block graph if it is chordal and diamond-free.
Removing an edge from a 4-clique $K_{4}$ results in a diamond. Diamond-free requires that no induced subgraph $G|_{R^{\prime}}$ is a diamond, where $R^{\prime}\subseteq R$ is any subset of vertices. We now present the main result relating block graphs to Berge-acyclic hypergraphs:
Lemma 31.
The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Together with the fact that every block graph is geodetic (it has a unique shortest path between any two vertices) [31], Section˜5.1 implies the following corollary.
Corollary 32.
Theorem 33.
Let $L$ be the line graph of a Berge-acyclic hypergraph $H$, and $\mathsf{P}(r,r^{\prime})$ be the shortest path in $L$ between $r,r^{\prime}\in R(L)$, $T_{r}=\bigcup_{r^{\prime}\in R(L)}\mathsf{P}(r,r^{\prime})$ is the unique canonical join tree for $H$ rooted at $r$.
@@ 5.1 Existence & Uniqueness of the Canonical Join Tree
A spanning tree of $L$ is its join tree by Section˜5. To show $T_{r}$ is a join tree, it is sufficient to show $T_{r}$ is spanning, connected and acyclic.
$T_{r}$ is spanning and connected because it contains a path from $r$ to each $r^{\prime}\in R(L)$.
We prove $T_{r}$ is acyclic by induction on the distance $\mathsf{dist}(r,r^{\prime})$ between $r$ and $r^{\prime}$. Let $R_{d}=\left\{r^{\prime}\in R(L)\mid\mathsf{dist}(r,r^{\prime})\leq d\right\}$. $R_{0}=\{r\}$ contains only the root. The subgraph $T_{r}|_{R_{0}}$ is trivially acyclic. Assuming that $T_{r}|_{R_{d>0}}$ is acyclic, we consider a vertex $r^{\prime}\in R_{d+1}\setminus R_{d}$. Section˜5.1 guarantees a unique shortest path between each pair of vertices $r,r^{\prime}\in R(L)$. Each $r^{\prime}$ is connected to a unique neighbor $r^{\prime\prime}\in R_{d}\setminus R_{d-1}$ that is at distance $d$ from $r$. Otherwise, there are at least two distinct shortest paths from $r$ to $r^{\prime}$. Therefore $T_{r}|_{R_{d+1}}$ is acyclic, and $T_{r}$ is a join tree.
The join tree $T_{r}$ is canonical, because the path from $r$ to each $r^{\prime}\in R(L)$ is the shortest and therefore minimizing the depth $0pt(T_{r},r^{\prime})$. The canonical tree is unique by Section˜5.1. ∎
5.2 Construction of the Canonical Join Tree
Given a Berge-acyclic hypergraph $H$ and $r\in R(H)$ as the root, we can construct the canonical tree $T_{r}(H)$ by running Algorithm˜1. We first present a useful lemma on the structure of the canonical join tree:
@@ 5.2 Construction of the Canonical Join Tree
We now show that the MCS algorithm constructs the canonical join tree.
Theorem 35.
Given a Berge-acyclic hypergraph $H$, Algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
@@
Algorithm˜1 labels each hyperedge in $R(H)$ in ascending order from the root $r_{1}=r$ to $r_{n=|R(H)|}$. When Algorithm˜1 has labeled $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{(r_{2},p(r_{2})),\ldots,(r_{i},p(r_{i}))\}$.
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, Algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜33. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in Figure˜5 of Appendix˜F, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in Figure˜5. By Section˜5.2, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases:
@@
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})=\chi(e_{r})$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi(e_{l})\neq\chi(e_{r})$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
@@
6 Converting a Binary Join Plan to a Join Tree
A recent approach [20, 2] that converts a binary join plan into a join tree has gained popularity as it allows system builders to leverage existing query optimizers designed for binary join plans. In this section, we focus on an algorithm by Hu et al. [20] that converts left-deep linear join plans into join trees. The algorithm can convert all acyclic queries in standard benchmarks, including those in Table˜1 of Appendix˜F. We prove that the algorithm converts any connected left-deep linear join plan into a join tree if and only if the query is $\gamma$-acyclic. Our result yields a new characterization of $\gamma$-acyclic queries. We formally define binary join plans and describe the algorithm by Hu et al. [20] with Algorithm˜6 in Appendix˜F.
Definition 36.
A left-deep linear plan is a sequence of relations $r_{1},r_{2},\ldots,r_{n}$. It is connected if for each $r_{i\geq 2}$, $\exists r_{j<i}:r_{i}\cap r_{j}\neq\emptyset$.
Given a left-deep linear plan, Hu et al. [20] generate a join tree with Algorithm˜6 in Appendix˜F. The algorithm chooses the first relation $r_{1}$ as the root and iterates through the rest of the plan. For each relation $r_{i\in[2,n]}$, it finds the first relation $r_{j}$ that contains all attributes shared by $r_{i}$ with all previous relations $r_{i}\cap\bigcup r_{k<i}$, and assigns $r_{j}$ as the parent of $r_{i}$. The algorithm constructs a join tree if it finds a parent for each $r_{i}$.
Theorem 37.
The “only if” direction is more difficult. We prove its contrapositive by constructing a $\gamma$-cycle from a query plan that is not the reverse of a GYO-reduction order. We start with a simple observation that directly follows from the definition of GYO-reduction order.
@@ 6 Converting a Binary Join Plan to a Join Tree
Given a pair of incomparable relations $\hat{r}$ and $\bar{r}$ with respect to $\mathring{r}$, we can find at least one variable in each relation that does not appear in the other relation but appears in $\mathring{r}$, namely $\exists\bar{x}\in\bar{r}\cap\mathring{r}:\bar{x}\notin\hat{r}$ and $\exists\hat{x}\in\hat{r}\cap\mathring{r}:\hat{x}\notin\bar{r}$. We say that $\hat{x}$ is a dangling variable of $\hat{r}$ and $\bar{x}$ is a dangling variable of $\bar{r}$. We are now ready to prove Theorem˜37. Our strategy is to derive a $\gamma$-cycle from any orphan relation, leading to a contradiction.
Proof of Theorem˜37.
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By Section˜6, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, Section˜6 guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.

### v3

@@ 11    $x\leftarrow\text{a vertex in }\chi(e)\setminus\chi(p(e))$
From $L(H)$ we first retrieve query hypergraph $H$. This is done in time linear in the size of the line graph. algorithm˜1 constructs an MCS tree $T$ in $\mathcal{O}(|H|)$. With the inputs $L(H)$ and $T$, we can run algorithm˜2 to construct $G^{\equiv}(H)$ in $\mathcal{O}(|L|)$ by Theorem˜26. Finally, letting $n$ be the number of vertices and $m$ the number of edges in $L(H)$, the spanning tree enumeration algorithm by Kapoor and Ramesh [24] on $G^{\equiv}(H)$ enumerates all join trees $|\mathcal{T}(H)|$ in $\mathcal{O}(m+n+|\mathcal{T}(H)|)=\mathcal{O}(|L(H)|+|\mathcal{T}(H)|)$.
C.2 Missing Proofs in section˜4.3
We prove Theorem˜28 below.
Theorem˜28 The join trees of a $\gamma$-acyclic $H$ can be enumerated in time $\mathcal{O}(|H|+|\mathcal{T}(H)|)$.
@@
By modifying the algorithm˜1 to algorithm˜3, where changes are highlighted in red, we can construct a weighted MCS tree $T$ and an unweighted line graph $L(H)$ in $\mathcal{O}(|H|+|E(L(H))|)$. We simplify algorithm˜2 to algorithm˜4 where the changes are highlighted in red. As $\gamma$-acyclicity guarantees that every non-tree edge $(r_{i},r_{j})$ is an MST edge, we check its LCA $l$ and the weights of LCA edges $\omega_{1},\omega_{2}$ to determine how to slide it. If the shallower endpoint is the LCA of the deeper $r_{i}=l$, we slide it to be parallel with the only LCA edge as in Line 4. Otherwise, if the LCA edges have different weights, we slide $\{r_{i},r_{j}\}$ to be parallel to the heavier one as in Line 4 and Line 4. Finally if they have the same weight, we slide $\{r_{i},r_{j}\}$ to form a triangle with them as in Line 4. Those four cases are exhaustive.
The number of join trees is lower bounded by the number of non-tree edges in its line graph, namely $|\mathcal{T}(H)|=\Omega(|E(L(H))|-|E(T)|)$. The total time complexity of enumerating all join trees is $\mathcal{O}(|H|+|E(L(H))|+|\mathcal{T}(H)|)=\mathcal{O}(|H|+|\mathcal{T}(H)|)$. ∎
Appendix D Missing Proofs in section˜5
Proof.
Fagin [15] showed that every Berge-acyclic hypergraph is linear. We prove the other direction by contradiction, assuming that $H$ is $\alpha$-acyclic and linear. We also assume that there is a Berge cycle $(r_{0},x_{0},\dots,r_{k-1},x_{k-1})$ of length $k\geq 2$. Without loss of generality, we suppose $r_{1}$ is the first hyperedge on the cycle to be removed by GYO reduction. Then $r_{1}$’s parent $r_{p}$ must satisfy $r_{p}\cap r_{1}\supseteq(r_{1}\cap\bigcup_{i\in[k]\wedge i\neq 1}r_{i})=\{x_{0},x_{1}\}$, contradicting $H$’s linearity. ∎
section˜5 The line graph $L$ of a Berge-acyclic hypergraph $H$ is a block graph.
Proof.
A Berge-acyclic hypergraph is $\gamma$-acyclic [15]. Zhu proves that the line graph of a $\gamma$-acyclic hypergraph is chordal [49]. We now prove that the line graph is diamond-free by showing that every diamond must be part of a $K_{4}$. Assuming a diamond $(r_{1},r_{2},r_{3},r_{4})$ in $L$ without the edge $(r_{1},r_{3})$, it can be treated as two triangles $(r_{1},r_{2},r_{4})$ and $(r_{2},r_{4},r_{3})$ sharing an edge $(r_{2},r_{4})$. We now consider the possible edge labels, $r_{i}\cap r_{j}$, in each triangle:
Case 1:
Case 2:
Since the two triangles share the edge $(r_{2},r_{4})$, all their edges are labeled with the same variable $x$, namely $x\in r_{1}\cap r_{2}\cap r_{3}\cap r_{4}$. There has to be an edge $(r_{1},r_{3})\in E(L)$, contradicting the assumption that $(r_{1},r_{2},r_{3},r_{4})$ is a diamond. ∎
@@ Appendix D Missing Proofs in section˜5
Let $r_{0}=r$, section˜5 requires that $|r_{i-1}\cap r_{i}|=1$ for all $i\in[k]$. By Theorem˜35, the path $r_{0},\ldots,r_{k}$ is the shortest path from $r_{0}$ to $r_{k}$. The variable shared by two adjacent vertices $r_{i-1},r_{i}$ is the edge label $r_{i-1}\cap r_{i}$. The edge labels along the path are all distinct. Otherwise, two consecutive edges with the same label allow for shortening the path by removing either, contradicting the shortest path property. Two non-consecutive edges with the same label violate the running intersection property. ∎
Theorem˜36 Given a Berge-acyclic hypergraph $H$, algorithm˜1 from $r\in R(H)$ constructs the canonical tree $T_{r}(H)$.
Proof.
Denoting the edges labeled by algorithm˜1 as $r=r_{1},\dots,r_{n=|R(H)|}$, when algorithm˜1 has labeled the first $i\in[n]$ hyperedges, we consider a graph $G_{i}$ whose vertices are the labeled hyperedges $\{r_{1},\ldots,r_{i}\}$ and edges are the parent-child relationships $\{\{r_{2},p(r_{2})\},\ldots,\{r_{i},p(r_{i})\}\}$.
To prove the claim by induction, we show a loop invariant that $G_{i}$ is a subtree of $T_{r}$ for all $i\in[n]$. An empty graph is trivially a subtree of any graph. When the root $r$ is labeled as $r_{1}$, this graph of a single vertex is also trivially a subtree of $T_{r}$. Assuming $G_{k}$ is a subtree of $T_{r}$, algorithm˜1 proceeds to label $r_{k+1}$ and assigns its parent $p(r_{k+1})=r_{p\leq k}$. Suppose for the sake of contradiction that $r_{k+1}$ has a different parent $r_{q}$ in $T_{r}$. Let us consider the possible relationships between $r_{p}$ and $r_{q}$. First, $r_{q}$ cannot be a descendant of $r_{p}$, otherwise there would be a shorter path from $r_{k+1}$ to the root $r$ going through $r_{p}$ instead of $r_{q}$, violating Theorem˜35. Suppose $r_{q}$ is an ancestor of $r_{p}$. By the running intersection property, $r_{q}\cap r_{k+1}$ must contain $r_{p}\cap r_{k+1}$, but that would imply every vertex in $r_{p}\cap r_{k+1}$ is already marked by $r_{q}$ and the algorithm would not have assigned $r_{p}$ as the parent of $r_{k+1}$. Therefore, $r_{p}$ and $r_{q}$ are not descendants of each other. Denote their lowest common ancestor in $T_{r}$ as $r_{a}=\mathsf{LCA}(r_{p},r_{q})$. Let $\lambda(r_{p},r_{q})=\{e_{l},e_{r}\}$ as shown by the solid lines in fig.˜7, $\mathsf{P}_{p}$ be the path from $r_{a}$ to $r_{p}$ excluding $e_{l}$ and $\mathsf{P}_{q}$ be the path from $r_{a}$ to $r_{q}$ excluding $e_{r}$ as shown by the dashed lines in fig.˜7.
Figure 7: Possible cases for how the relation $r_{k+1}$ attaches to the partial join tree $G_{k}$ during the construction of the canonical join tree by Maximum Cardinality Search.
By appendix˜D, the variables on $\mathsf{P}_{p}$ and $\mathsf{P}_{q}$ are all distinct. Consider the following possible cases where $\chi^{l}=\bigcap_{r\in\rho(e_{l})}{r}$ and $\chi^{r}=\bigcap_{r\in\rho(e_{r})}{r}$:
@@
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi^{l}=\chi^{r}$, then $\mathsf{P}_{p},\mathsf{P}_{q}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}=r_{k+1}\cap r_{q}$ and $\chi^{l}\neq\chi^{r}$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi^{l}=\chi^{r}$, then $\mathsf{P}_{p},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
@@
If $r_{k+1}\cap r_{p}\neq r_{k+1}\cap r_{q}$ and $\chi^{l}\neq\chi^{r}$, then $\mathsf{P}_{p},r_{a},\mathsf{P}_{q},r_{k+1}$ form a Berge-cycle.
@@
Appendix E Missing Algorithm and Proofs in section˜6
Theorem˜38 A query is $\gamma$-acyclic if and only if every connected left-deep linear join plan for the query is the reverse of a GYO-reduction order.
We first prove “only if” direction of Theorem˜38 by contradiction, assuming a query plan with a $\gamma$-cycle that is not the reverse of a GYO-reduction order.
@@ Appendix E Missing Algorithm and Proofs in section˜6
Given a pair of incomparable relations $\hat{r}$ and $\bar{r}$ with respect to $\mathring{r}$, we can find at least one variable in each relation that does not appear in the other relation but appears in $\mathring{r}$, namely $\exists\bar{x}\in\bar{r}\cap\mathring{r}:\bar{x}\notin\hat{r}$ and $\exists\hat{x}\in\hat{r}\cap\mathring{r}:\hat{x}\notin\bar{r}$. We say that $\hat{x}$ is a dangling variable of $\hat{r}$ and $\bar{x}$ is a dangling variable of $\bar{r}$. We are now ready to prove Theorem˜38. Our strategy is to derive a $\gamma$-cycle from any orphan relation, leading to a contradiction.
Proof of Theorem˜38.
To prove the “only if” direction by contradiction, we assume that a connected left-deep linear plan for a $\gamma$-acyclic query is not the reverse of a GYO-reduction order. By appendix˜E, there is an orphan $\mathring{r}$ in the plan. Assuming that $\mathring{r}$ is the $t$-th relation in the plan, we now consider the prefix of the plan up to $\mathring{r}$, namely $\mathsf{pre}({t})=r_{1},\ldots,r_{t-1},\mathring{r}$. Let $\hat{r}\in\mathsf{pre}({t-1})$ be a maximal relation with respect to $\geq_{\mathring{r}}$, appendix˜E guarantees an $\bar{r}\in\mathsf{pre}({t-1})$ incomparable with $\hat{r}$. There are at least two dangling variables $\bar{x}\in\bar{r}\cap\mathring{r}\setminus\hat{r}$ and $\hat{x}\in\hat{r}\cap\mathring{r}\setminus\bar{r}$.